UCL Ear Institute


3 March 2010 minutes

UCL Ear Institute, Health & Safety Meeting Minutes


  • Professor David McAlpine (DM)
  • Graham Nevill (GN)
  • Adam Pyne (AP)
  • Sally Dawson (SD)
  • Mr John Kelly (JK)
  • Ms Miriam Gomez (MG)
  • Nico Daudet (ND)
  • Cheryl Overington (CO)

 1. Apologies for Absence

Apologies were received from: Bill Potter (BP), Edwina Turner (ET) (minute secretary)

2. Minutes of the previous meeting

Agreed the minutes must be proof read and approved prior to circulation. In future, ET to produce the Minutes and GN to amend and approve them.

3. Matters Arising

(a) New safety positions within the EI

(b) GM Officer – Nico Daudet to take on position instead of Bjorn Christianson as he has more active involvement in this work.

Service contracts: Noted service contracts across CAR are due for renewal. CO estimated that there would be £5000 approx available towards the service contract costs (exact figure to be confirmed but if less could be made up to £5k).

Fire Marshalls: DRUK need a FEM for the 2nd floor, Rada confirmed this was in hand but the person needed to attend the course. FEM also needed for 2nd floor of teaching area and agreed Cherilee would be asked to do this as her position is full-time. Noted Bill Potter also to be a FEM.

4. Risk Assessments

GN reported that all lab risk assessments had now been completed. CO noted that offices may still need to be completed and she will review this.

5. Departmental Inspection Reports
Monthly inspection topics are protective equipment and UV safety. In the case of the former, there are adequate and suitable types for lab work currently done and staff are aware of the need for their use. In terms of UV Safety a sign had been placed in the relevant area alerting people to the potential hazard. Lab coats – currently no way to arrange for them to be cleaned and it was discussed whether it would be possible for this to be organised through the hospital. AP to discuss with Sunday in the hospital and with Sam at UCL.

Noted by GN that the fire evacuation test showed there to be a problem with one external door which was stiff to open. AP thinks it is resolved but will check and action if necessary.

Graham Nevill and Miriam Gomez undertook a Dept Inspection as required on 2 March. Findings:

2nd Floor

An excessive amount of waste and clutter is in the lab area, mainly packaging which represents a fire hazard. DM requested a major clear-out blitz by all staff. Noted the skip would be available for a few more days.

Temperature – lab far too hot resulting in doors being propped open. AP reported that this had been discussed at a meeting the previous day and hoped that he might be able to fix but needed the information from Andy Forge. Also E&F (Tony Overbury) is trying to resolve the problem. DM requested that all maintenance issues and work go through AP (Andy Forge or GN in AP’s absence). Noted the filters were changed in the summer and may not have been completed properly. Also noted the first floor is freezing cold. AP to report the problems again + e-mail Tony Overbury and cc Brian Kavanagh noting it is a safety issue.

1st Floor, Tissue Culture

A number of ceiling lights not working but may have been deliberately inactivated – need for signs asking people not to put on the lights. AP to check the lights on Saturday to see if they are working and if not to fix them and put up signs.

Noted there is also a problem with the lights on the second floor. AP will also check these on Saturday but has recently replaced bulbs and thinks there might be a problem with the lights themselves.

Ground floor

Noted switch in office area hanging out of wall. AP to fix on Saturday but noting he had tried previously and believe it to have been fixed.

Noted door to open office area does not shut properly. GN reported he had lubricated it and AP had changed the speed with little improvement. AP thought people were too hasty and didn’t wait for the mechanism to release which is causing problems with the lock. AP to look again to see if further improvement is possible.


Noted clutter in fire escapes. AP reported that this is a result of the building works in B34 and B35 and that UCL E&F will collect soon once the work has been completed. AC to clear out the storage space. Noted space needed for the laminator but agreed that possibly the laminator should be disposed of and a small one purchased suitable for signage etc but not full size posters.

Back room – being used by Joerg Albert and Simon Gain. Noted that blinds might be required.

Noted a lot of junk in B17 (and on the 1st floor). Church door area - filing cabinets and clutter including bags for shredding. Noted this is a fire exit area for Lecture Theatre and students and therefore vital there is clear access. Shredding awaiting collection by UCL. Minal Shah had been using the filing cabinets but had been in the process of emptying them prior to leaving.

B17, microscopy and corridor area: Only 1 light working. First Aid box needs mounting on the wall by sink with appropriate signage. AP to action.

6. DRUK Reports

Nothing to report.

7. Any Other Business

Major Incident Plan – some discussion around this subject but noted a separate meeting is scheduled to discuss this and take forward.

John Kelly raised the matter of screening for students working in the animal unit and if this is required. Noted that whilst screening forms part of the appointment process for staff, the situation with respect to students is unclear. CO to e-mail Occupational Health and seek advice.

Noted Fire Evacuation Marshalls are required to go on refresher courses. GN had already requested FEMs to do so but will send out another reminder.

8. Dates of next meeting

The date of the next meeting will be in June, but no specific date has been set.